OpenStudy (ashy98):

Create a compound inequality to represent the air pressure range. Can the air pressure in this town be 24 inches? Justify your answer by solving the inequalities in Part A. The average air pressure in another town is 29 inches, but the actual pressure is within 4 inches of the average. Write and solve an inequality to find the range of air pressure in this town.

OpenStudy (anonymous):

4/29=x/24 solve it now

OpenStudy (ashy98):

x=96/29~3.3103?

OpenStudy (ashy98):

@TheCoque???